OpenStudy (aperez):

write an expression for each phrase..........one half of the quotient of b and 4

jimthompson5910 (jim_thompson5910):

the quotient of b and 4 means b/4

jimthompson5910 (jim_thompson5910):

So what is "one half of the quotient of b and 4"

OpenStudy (aperez):

1/2 perhaps

jimthompson5910 (jim_thompson5910):

so multiply 1/2 by b/4 to get ???

OpenStudy (aperez):

2

jimthompson5910 (jim_thompson5910):

(1/2)*(b/4) = (1*b)/(2*4) = ?

OpenStudy (aperez):

sorry 8

jimthompson5910 (jim_thompson5910):

so the final answer is b/8

OpenStudy (aperez):

Thanks

jimthompson5910 (jim_thompson5910):

np